# Quantum (quadratic) gravity: replacing the massive tensor ghost with an inverted harmonic oscillator-like instability

### 07 Mar 2026 v1

@@ 5.2 Quantization of IHO spin–2 field with geometric superselection sectors
where $\pi_{s}=\dot{v}_{s}$. This structure differs from a ghost: the Hamiltonian is not negative definite. Rather, it possesses a hyperbolic quadratic form in phase space. The defining feature of the dual IHO is the negative sign in the kinetic term, while the spatial gradient and mass terms remain positive. The field equation reads
@@
This is the dual IHO: the instability can be seen easily in momentum space rather than configuration space. The energy shells satisfy
@@
The minimal $\mathcal{P}\mathcal{T}$-symmetric physical construction is therefore a direct-sum of conjugate sectors. Therefore, we quantize each transverse–traceless polarization of the massive spin–2 sector as a dual IHO. There is no global Fock vacuum that exists, no positive-frequency split can be defined, and no additional asymptotic particle states arise as we discussed earlier. Furthermore, we cannot interpret the IHO spin–2 mode as a particle, just as quantum fields in curved spacetime cannot have a particle interpretation. This exploits the elegance of QFTCS in shedding light on the nature of quantum gravity. In both physical contexts, the healthy instability, like IHO, plays a universal role. In the context of QFTCS, DQFT splits a quantum state as a direct-sum of 4-components with pair of parity conjugate regions inside and outside the horizon. Note that the concept of time is different from the interior to exterior, therefore, they form geometric superselection sectors and a single quantum state is projected as "mirror" images on either side of the gravitational horizons. It is exactly this construction shown to save unitarity and resolve information paradoxes in curved spacetime [97, 96, 95].
@@ 6 A safe beginning of the Universe with the unitary quadratic quantum gravity
What remains is a restricted class of early-time behaviors in which anisotropy is dynamically regulated so that $a^{3}W^{2}$ remains integrable. In this regime, the additional spin–2 sector of quadratic gravity plays a crucial structural role. For $\beta>0$, the massive spin–2 mode is dynamically a dual IHO and does have any asymptotic particle excitations, it does not enlarge the physical spectrum. Instead, it acts as a non-particle instability channel in the UV regime, redistributing anisotropic energy and suppressing the shear growth that would otherwise drive $W^{2}$ and the quadratic action to divergence.
@@ 6.1 Starobinsky inflation as an emergent scenario
where the oscillatory factors later result in the particle production by the decay of the scalaron, which renders the Universe into the radiation phase [129, 80]. It is important to note that the quadratic gravity is renormalizable, even with couplings to matter fields [39, 36]. However, for the UV renormalizability, the Higgs field must be coupled non-minimally to quadratic gravity [36]. Coupling of matter fields to quadratic gravity can be written as
@@ 7.1.1 Scalar perturbations:
where $|0\rangle_{\rm qdS}$ is the quasi-de Sitter vacuum where we quantize the MS variable $v$. In (7.12), the factor $\frac{1}{2a^{2}Q_{s}}$ comes from rescaling the field operator $\hat{v}$ by the classical factor $z_{s}$. The quantization procedure we follow is unitary QFTCS that relies on DQFT [93, 63, 64, 65], which we present in the next section. Assuming the Bunch-Davies vacuum solution for the mode function (in de Sitter approaxmimation)
@@
which is exactly the same as $R+R^{2}$ gravity [48] since the Weyl square term does not contribute to the scalar sector at the linearized level. Evaluating quantities at horizon crossing $k=aH$, we can relate derivatives with respect to $k$ and the number of e-folds $N$ (from the end of inflation) via
@@ 7.1.2 Tensor perturbations
where we applied the approximation $F\approx 24\alpha H^{2}$. The result in (7.39) matches with the derivations in [49, 10, 121] From the above tensor-to-scalar ratio expression, we can deduce that $12\alpha>\beta>0$. For the scalaron mass $m_{0}\sim 1.3\times 10^{-5}M_{p}$ (taking $N=55$), we get

### 25 Mar 2026 v2

@@ 5.1 Quantization of IHO spin–2 field with geometric superselection sectors
where $\pi_{s}=\dot{v}_{s}$. This structure differs from a ghost: the Hamiltonian is not negative definite, rather indefinite. Rather, it possesses a hyperbolic quadratic form in phase space. The defining feature of the dual-IHO is the negative sign in the kinetic term, while the spatial gradient and mass terms remain positive. The field equation reads
@@
This is the dual-IHO: the instability can be seen easily in momentum space rather than configuration space. The energy shells satisfy
@@
The minimal $\mathcal{P}\mathcal{T}$-symmetric physical construction is therefore a direct-sum of conjugate sectors. Therefore, we quantize each transverse–traceless polarization of the massive spin–2 sector as a dual-IHO. There is no global Fock vacuum that exists, no positive-frequency split can be defined, and no additional asymptotic particle states arise as we discussed earlier. Furthermore, we cannot interpret the dual-IHO spin–2 mode as a particle, just as quantum fields in curved spacetime cannot have a particle interpretation. This exploits the elegance of QFTCS in shedding light on the nature of quantum gravity. In both physical contexts, the healthy instability, like IHO, plays a universal role. In the context of QFTCS, DQFT splits a quantum state as a direct-sum of 4-components with pair of parity conjugate regions inside and outside the horizon. Note that the concept of time is different from the interior to exterior, therefore, they form geometric superselection sectors and a single quantum state is projected as "mirror" images on either side of the gravitational horizons. It is exactly this construction shown to save unitarity and resolve information paradoxes in curved spacetime [103, 102, 101].
5.1.1 Källén–Lehmann representation and the principal value propagator of the dual-IHO spin-2 field
Here we focus on understanding the propagator of the dual-IHO spin–2 sector which is fundamentally different from the Feynman propagator in the standard QFT.
According to the Källén–Lehmann (KL) representation [84, 107, 135], the full non-perturbative Feynman propagator of a scalar field (with $i\epsilon$ prescription) is
$\tilde{\Delta}(k^{2})=\frac{-i}{k^{2}+m^{2}-i\epsilon}+\int_{0}^{\infty}ds\,\frac{-i\rho(s)}{k^{2}+s-i\epsilon}$ (5.15)
The spectral density is defined as:
$\rho(s)\equiv\sum_{n}|\langle k,n|\varphi(0)|0\rangle|^{2}\,\delta(s-M_{n}^{2})$ (5.16)
This packages all the multiparticle matrix elements into a single function of $s\geq 4m^{2}$. The $M_{n}\geq 2m$ here is the invariant mass of the multiparticle states. It satisfies $\rho(s)\geq 0$ and $\rho(s)=0$ for $s<4m^{2}$.
Here $|0\rangle$ is the vacuum state and $|n\rangle$ runs over all physical one-particle states. Three properties of (5.15)–(5.16) are essential for what follows.
  1. 1.
    The spectral domain is timelike. The integration variable $m^{2}\geq 0$ corresponds to $k^{2}=-m^{2}\leq 0$, i.e. to timelike or lightlike momenta. A propagator pole at spacelike momentum $k^{2}=\mu^{2}>0$ corresponds to $m^{2}=-\mu^{2}<0$, which lies outside the domain of integration in (5.15). No standard Feynman propagator with a spacelike pole can be represented in the form (5.15).
  2. 2.
    $\rho\geq 0$ requires a normalizable vacuum. The matrix element $\langle 0|\phi(0)|n\rangle$ is only defined if the bra $\langle 0|$ exists as a normalizable state, $\langle 0|0\rangle=1$. If no such state exists, the spectral density (5.16) is undefined; one must set $\rho=0$.
  3. 3.
    The $\delta$-function piece requires a normalizable one-particle state. Each term in (5.16) contributes a delta function $\delta(s-M_{n}^{2})$ to $\rho$ only if the state $|n\rangle$ is a normalizable eigenstate with a definite mass $M_{n}$. Non-normalizable (Dirac-delta normalised) scattering states contribute a continuous measure to $\rho$, not isolated delta functions; but even this continuous contribution requires a normalizable vacuum $|0\rangle$ for the matrix element to be defined.
We have argued in Sec. 2 that the dual-IHO ground state is not in $L^{2}(\mathbb{R})$. This result holds regardless of whether additional quantization conditions are imposed. Those conditions select a discrete set of allowed energies but they do not alter the asymptotic behaviour of the eigenfunctions within each sector. The wave functions at the discrete energies are still parabolic cylinder functions, which diverge and are still not in $L^{2}(\mathbb{R})$.
The KL spectral density of the dual-IHO spin-2 field vanishes for two independent reasons. First, there is no normalizable vacuum. Since no eigenfunction of $H_{s,\lambda}$ is in $L^{2}(\mathbb{R})$, there is no normalizable state $|0\rangle$ with $\langle 0|0\rangle=1$. The matrix element $\langle 0|\phi(0)|n\rangle$ in (5.16) is undefined. The spectral density is therefore undefined and must be set to zero: $\rho(\mu_{2}^{2})=0$. Second, the pole lies outside the KL spectral domain. The dual-IHO propagator has a pole at $k^{2}=\mu_{2}^{2}>0$ (spacelike). In the KL representation (5.15) this would correspond to $s=-\mu_{2}^{2}<0$, which lies outside the integration domain $m^{2}\geq 0$. No physical particle with real mass can produce a spacelike pole. Even if one attempted to force a spectral contribution at $s=-\mu_{2}^{2}$, it would violate the positivity of the spectral domain and is inadmissible.
Both reasons independently enforce $\rho(\mu_{2}^{2})=0$. Together they establish that the dual-IHO spin-2 field carries no spectral weight at $k^{2}=\mu_{2}^{2}$. With $\rho(\mu_{2}^{2})=0$, the $\delta(k^{2}-\mu_{2}^{2})$ piece is absent from the propagator, not suppressed by kinematics, not removed by a prescription, but identically absent as a consequence of the Hilbert space structure. The propagator retains only its dispersive (real) part:
$\Delta_{\rm Dual-IHO}(k)\;=\;\mathrm{PV}\frac{i}{k^{2}-\mu_{2}^{2}}$ (5.17)
This is the unique prescription consistent with $\rho=0$: it has no imaginary part, no absorptive contribution, and no on-shell piece.
The distinction from plane waves is worth emphasising. Free field plane wave states $|{\bf k}\rangle=a_{\bf k}^{\dagger}|0\rangle$ are not in $L^{2}(\mathbb{R})$, they satisfy $\langle{\bf k}|{\bf k}^{\prime}\rangle=\delta^{(3)}({\bf k}-{\bf k}^{\prime})$, yet the free scalar field has a perfectly well-defined Feynman propagator. The reason is that the vacuum $|0\rangle$ is normalizable: $\langle 0|0\rangle=1$, so the KL matrix element $\langle 0|\phi(0)|{\bf k}\rangle$ is defined. For the dual-IHO there is no normalizable state to place in the bra position regardless of the excited states. The decisive condition is the non normalizability of the ground state, not of the excited states.
In the work [82] it was proved that any Lorentz invariant interacting QFT containing a tachyonic (or IHO) field violates perturbative unitarity beyond tree level. Their proof proceeds by decomposing the cut tachyon propagator as
$\widetilde{\Delta}_{\rm cut}(k)\;=\;\underbrace{2\pi\,\theta(k^{0})\,\delta(k^{2}-\mu^{2})}_{\Delta_{A}}\;+\;\underbrace{\frac{k^{0}\,\theta(\mu^{2}-k^{2})}{(k^{2}-\mu^{2})\sqrt{\mu^{2}-k^{2}}}}_{\Delta_{B}}\,$, (5.18)
and showing that the $\Delta_{B}$ terms accumulate non-cancelling imaginary contributions at loop level that cannot be written as a sum over physical intermediate states, violating the optical theorem. Thus, the $\Delta_{B}$ contributions generate imaginary parts in $\mathrm{Im}(\mathcal{M}_{ii})$ that do not correspond to any physical on-shell intermediate state and therefore cannot appear on the right-hand side as $|\mathcal{M}_{if}|^{2}$ for any state $f$. The equality is therefore broken:
$2\,\mathrm{Im}(\mathcal{M}_{ii})=\underbrace{\sum_{f}\int d\Pi_{f}\,|\mathcal{M}_{if}|^{2}}_{\text{physical cuts}}+\underbrace{\Delta_{B}\text{-contributions}}_{\text{no physical state}}$ (5.19)
The left-hand side contains extra imaginary parts with no physical interpretation on the right-hand side, and the optical theorem, which is the diagrammatic statement of unitarity, is violated. The principal value prescription removes $\Delta_{B}$ exactly, restoring the equality.
First of all it is important to notice one structural difference between tachyonic (IHO) and dual-IHO field. Hamiltonian density of an IHO (tachyonic) field reads as
$\mathcal{H}_{T}=\frac{1}{2}\pi_{t}^{2}+\frac{1}{2}(\nabla v_{t})^{2}-\frac{1}{2}\mu^{2}v_{t}^{2},$ (5.20)
Though (5.20) indefinite like (5.4), for $k^{2}>\mu^{2}$, we could still define a normalizable states since the solutions are oscillatory [82]. For the dual-IHO field, we cannot have normalizabile states for any $k$.
Both $\Delta_{A}$ and $\Delta_{B}$ in (5.18) require $\rho(s)\neq 0$, they are the on-shell and imaginary pole pieces of a propagator with spectral weight at $k^{2}=\mu^{2}$. Since $\rho(s)=0$ for the dual-IHO spin–2 field, both pieces are absent:
$\Delta_{A}^{\rm IHO}\;=\;0\,,\qquad\Delta_{B}^{\rm IHO}\;=\;0\,.$ (5.21)
The aforementioned proof requires a Feynman propagator $i/(k^{2}-\mu^{2}-i\varepsilon)$ for the tachyonic (IHO) mode. This in turn requires a normalizable vacuum from which to define the KL spectral density. The dual-IHO satisfies neither condition. So the analysis of [82] therefore does not apply. It follows that its hypothesis, the existence of a tachyonic field with nonzero spectral density, is not satisfied.
The PV propagator (5.17) has zero imaginary part at all momenta, including the region $|\mathbf{k}|<\mu_{2}$ where the imaginary poles of the Feynman propagator would reside. At every loop order, the Cutkosky cutting rule applied to $\Delta_{\rm dual-IHO}$ gives
${\rm Cut}\,\Delta_{\rm Dual-IHO}(k)\;=\;\tfrac{1}{2}\bigl[2\pi\delta(k^{2}-\mu_{2}^{2})\theta(k^{0})\bigr]+\tfrac{1}{2}\bigl[-2\pi\delta(k^{2}-\mu_{2}^{2})\theta(k^{0})\bigr]\;=\;0\,$, (5.22)
so the dual-IHO spin-2 mode contributes zero imaginary part to any loop amplitude at any loop order. The optical theorem
$2\,\mathrm{Im}\,\mathcal{M}(a\to a)\;=\;\sum_{f}\int d\Pi_{f}\,|\mathcal{M}(a\to f)|^{2}$ (5.23)
is satisfied with the sum running only over physical asymptotic states, the massless graviton and the scalaron. The dual-IHO spin-2 contributes to neither side.
In summary, the non-normalizability of the dual-IHO ground state, derived from the parabolic cylinder function analysis of Eq. (2.20), forces the spectral density to vanish, $\rho(s)=0$, in the KL representation. This uniquely fixes the propagator to its principal value form in Eq. (5.17), eliminating any discontinuity across physical cuts at all loop orders. Consequently, the premises of the analysis presented in [82] are not satisfied, and unitarity is preserved. The entire chain is dictated by the structure of the theory itself, without the need for ad hoc prescriptions.
5.2 The dual-IHO spin–2 field, unitarity cuts and renormalization
Thus, the spin–2 sector has the schematic form (in the DQFT)
$D^{\pm(2)}_{\mu\nu\rho\sigma}(k)=\frac{\mp i}{M_{p}^{2}}\,P^{(2)}_{\mu\nu\rho\sigma}\left[\frac{1}{k^{2}\mp i\epsilon}-\mathrm{PV}\!\left(\frac{1}{k^{2}-\mu_{2}^{2}}\right)\right]\,$, (5.25)
where $\pm$ correspond to the parity conjugate geometric superselection sectors with opposite arrows of time (See Sec. 3). Notice that the dual-IHO spin–2 no $\delta(k^{2}-\mu_{2}^{2})$ support from the extra spin–2 pole. This implies, in quadratic gravity with a spin–2 dual-IHO pole at $k^{2}=\mu_{2}^{2}>0$, it does not contribute to the optical theorem at tree-level for physical graviton scattering since it cannot appear as an on-shell intermediate state in any unitarity cut (See panel (a) in Fig. 4). To be more precise, For $2\to 2$ a (forward) elastic scattering of physical massless external gravitons with momenta $p_{1},p_{2}\to p_{3},p_{4}$ obeying $p_{i}^{2}=0$ and $p_{i}^{0}>0$, one has the kinematic inequalities
$s\leq 0,\qquad t\geq 0,\qquad u\geq 0,$ (5.26)
with strict $s<0$ in the center-of-mass frame (since $k_{s}=(E_{1}+E_{2},\mathbf{0})$ implies $s=k_{s}^{2}=-(E_{1}+E_{2})^{2}$). Although the momentum transfers $k_{t},k_{u}$ are spacelike ($t,u\geq 0$), the optical theorem for forward $2\to 2$ scattering is controlled by the cut momentum flowing through the intermediate state, which is the total incoming momentum $k=k_{s}=p_{1}+p_{2}$. Therefore the on-shell condition for the dual-IHO pole, $k^{2}=\mu_{2}^{2}>0$, cannot be satisfied in the physical forward region. Therefore, the dual-IHO spin–2 propagator yields no contribution to $2\,\mathrm{Im}(M)$ in the forward unitarity cut. Since the optical theorem is precisely the statement that $2\,\mathrm{Im}(M)$ equals the sum over cut (on-shell) intermediate states,it follows that the IHO mode does not appear in that sum. Nevertheless, this spin–2 mode can still contribute off shell to tree amplitudes and to loop renormalization (real parts and UV divergences) through the principal value pieces of propagators, but it cannot contribute to the unitarity (optical theorem) sum because it never goes on shell in physical kinematics and it also its propagator must be defined by the principal value as we learned in the previous section.
$i\Pi(P)\;=\;\int\!\frac{d^{4}k}{(2\pi)^{4}}\,\mathcal{N}(k,P){\rm PV}\Big[\frac{i}{\big(k^{2}-\mu_{2}^{2}\big)}\Big]{\rm PV}\Big[\frac{i}{\big((P-k)^{2}-\mu_{2}^{2}\big)}\Big]$ (5.27)
where $\mathcal{N}(k,\,P)$ here contains the vertex factors. The location of branch points and cuts is controlled entirely by the Landau (pinch) conditions, i.e. by the possibility of putting the internal lines simultaneously on shell. For the present bubble these are
$k^{2}=\mu_{2}^{2},\qquad(P-k)^{2}=\mu_{2}^{2}.$ (5.28)
Subtracting the two equations in (5.28) gives
$k^{\mu}=a\,P^{\mu}+k_{\perp}^{\mu},\qquad P\!\cdot\!k_{\perp}=0.$ (5.30)
Then (5.29) fixes $a=\tfrac{1}{2}$, i.e.
$k^{\mu}=\frac{1}{2}P^{\mu}+k_{\perp}^{\mu},\qquad P\!\cdot\!k_{\perp}=0.$ (5.31)
Imposing the on-shell condition $k^{2}=\mu_{2}^{2}$ and using (5.31) yields
$\mu_{2}^{2}=k^{2}=\left(\frac{P}{2}\right)^{2}+k_{\perp}^{2}=\frac{P^{2}}{4}+k_{\perp}^{2}\quad\Longrightarrow\quad k_{\perp}^{2}=\mu_{2}^{2}-\frac{P^{2}}{4}.$ (5.32)
$P^{2}_{\ast}=4\mu_{2}^{2}\,$}. (5.33)
$\mathcal{M}^{\pm}_{1\text{-loop}}(p_{i})\;\sim\;\int\!\frac{d^{4}q}{(2\pi)^{4}}\;\frac{N(q,p_{i})}{\prod_{a}\big(D_{a}(q,p_{i})\mp i\epsilon\big)}.$ (5.34)
Here $p_{i}$ are external momenta, and:
  1. 1.
  2. 2.
To isolate the effect of a single dual-IHO spin–2 line, it is convenient to rewrite (5.34) as
$\mathcal{M}^{\pm({\rm dual-IHO)}}_{1\text{-loop}}\;\sim\;\int\!\frac{d^{4}q}{(2\pi)^{4}}\;F(q;p_{i})\;{\rm PV}\Big[\frac{1}{q^{2}-\mu_{2}^{2}}\Big],$ (5.35)
Here, $F(q;p_{i})$ is simply all the remaining in the integrand after factoring out the IHO-like spin–2 in the denominator. For fixed external kinematics, $F(q;p_{i})$ is a (typically smooth) function of $q$ away from the poles of the other denominators. Now, we proceed to identify the exact principal value and $\delta$ decomposition at the level of the loop.
Let us now verify that, in the UV limit, the principal value prescription preserves the renormalizability of quadratic gravity. For $|k^{2}|\gg\mu_{2}^{2}$ one finds
Therefore, at large loop momentum, the dual-IHO line falls off as $1/k^{2}$ (up to subleading powers), just like an ordinary propagator. Consequently, the UV degree of divergence of (5.35) is controlled by the same power counting as in the standard Feynman prescription; the IHO mode still contributes to UV divergences, counterterms, and running couplings.
which is supported only on the (finite) hypersurface $k^{2}=\mu_{2}^{2}$ in momentum space. This on-shell term generates the discontinuities (absorptive contributions) associated with cuts through that line, but it does not control the $|k|\to\infty$ region that produces UV divergences. Hence the dual-IHO prescription removes the on-shell (cut-producing) part of the dual-IHO line while retaining its off-shell contribution to renormalization and real (dispersive) loop corrections. Consequently, the dual-IHO mode can and does contribute virtually through the principal value part of its propagator. This yields:
  1. 1.
  2. 2.
    $D^{(2)}(k)\;\propto\;\left(\frac{1}{k^{2}}-{\rm PV}\left[\frac{1}{k^{2}-\mu_{2}^{2}}\right]\right)\,$, (5.38)
In summary, the dual-IHO prescription enforces a clean split between what is allowed and what is forbidden. The IHO-like spin–2 part is off-shell as there is no $\delta$-function support, thus its propagator is understood as a principal value, so loop integrals generate only real contributions, exactly the pieces needed for renormalization (counterterms and the real parts of self-energies/vertices). Since the $\delta$-supported part is what normally produces an absorptive contribution and hence a unitarity cut (via the optical theorem), the corresponding physical channel involving the spin–2 dual-IHO never opens. The dual-IHO spin–2 mode is therefore present as a virtual, renormalizing sector but is excluded from the spectrum of on-shell intermediate states. We also provided a further justification for this using the spectral representation of the dual-IHO propagator.
[…]
$D^{(2)}(k)$
(b) UV improvement via $1/k^{4}$ falloff.
Figure 4: Dual-IHO spin–2: absent from physical unitarity cuts (top), but improves UV behavior via $1/k^{4}$ falloff (bottom).
@@ 6 A safe beginning of the Universe with the unitary quadratic quantum gravity
What remains is a restricted class of early-time behaviors in which anisotropy is dynamically regulated so that $a^{3}W^{2}$ remains integrable. In this regime, the additional spin–2 sector of quadratic gravity plays a crucial structural role. For $\beta>0$, the massive spin–2 mode is dynamically a dual-IHO and does have any asymptotic particle excitations, it does not enlarge the physical spectrum. Instead, it acts as a non-particle instability channel in the UV regime, redistributing anisotropic energy and suppressing the shear growth that would otherwise drive $W^{2}$ and the quadratic action to divergence.
@@ 6.1 Starobinsky inflation as an emergent scenario
where the oscillatory factors later result in the particle production by the decay of the scalaron, which renders the Universe into the radiation phase [137, 83]. It is important to note that the quadratic gravity is renormalizable, even with couplings to matter fields [41, 38, 37]. However, for the UV renormalizability, the Higgs field must be coupled non-minimally to quadratic gravity [38, 37]. Coupling of matter fields to quadratic gravity can be written as
@@ 7.1.1 Scalar perturbations:
where $|0\rangle_{\rm qdS}$ is the quasi-de Sitter vacuum where we quantize the MS variable $v$. In (7.12), the factor $\frac{1}{2a^{2}Q_{s}}$ comes from rescaling the field operator $\hat{v}$ by the classical factor $z_{s}$. The quantization procedure we follow is unitary QFTCS that relies on DQFT [99, 65, 66, 67], which we present in the next section. Assuming the Bunch-Davies vacuum solution for the mode function (in de Sitter approximation)
@@
which is exactly the same as $R+R^{2}$ gravity [50, 121] since the Weyl square term does not contribute to the scalar sector at the linearized level. Evaluating quantities at horizon crossing $k=aH$, we can relate derivatives with respect to $k$ and the number of e-folds $N$ (from the end of inflation) via
@@ 7.1.2 Tensor perturbations
where we applied the approximation $F\approx 24\alpha H^{2}$. Structurally the result in (7.39) matches with the derivations in [51, 10, 129, 34] when applied to our case of $\beta>0$. From the above tensor-to-scalar ratio expression, we can deduce that $12\alpha>\beta>0$. For the scalaron mass $m_{0}\sim 1.3\times 10^{-5}M_{p}$ (taking $N=55$), we get